A certain type of battery has a mean shelf life of 750 with a standard deviation 0f 33 days

a. What is the probability that the shelf life of batteryis below 350 days

b. what is the probability that the shelf life of battery is between 717 and 816 days


1
Expert's answer
2022-03-11T01:41:15-0500

"\mu=750,\sigma=33,Z=(x-μ)/σ"

a. "P(x<350)=P(Z<(350-750)\/33)=P(Z<-12.12)"

From z distribution tables

"P(Z<-12.12)=0.000"

b. "P(717<x<816)=P((717-750)\/33<Z<(816-750)\/33)"

"=P(-1<Z<2)=P(Z<2)-P(Z<-1)"

"=0.9772-0.1587"

"=0.8185"